The hypotenuse of a right triangle is 22 m long. The
length of one leg is 10 m less than the other. Find
the lengths of the legs.

Found 2 solutions by mananth, nyc_function:
Answer by mananth(16946)

The hypotenuse of a right triangle is 22 m long. The
length of one leg is 10 m less than the other. Find
the lengths of the legs.
let one leg be x
the other leg will be x-10
x^2 +(x-10)^2=22^2
x^2+x^2-20x+100=484
2x^2-20x-384=0
2(x^2-10x-192)=0
x^2-10x-192=0
x1= 10+sqrt(100+768) / 2
x1=19.731
.
x2= 10-sqrt(100+768) / 2
x2=-9.731
Since x1 is positve x= 19.731
The other leg = 9.731
Checking
(19.731)^2 + (9.731)^2= 484 .00




Answer by nyc_function(2741)

(leg)^2 + (leg)^2 = (hypotenuse)^2
(x)^2 + (x - 10)^2 = (22)^2
This is all you need to find the length of each leg.
Can you do it now?